# An Extension of the Localized Artificial Diffusivity Method for Immiscible and High Density Ratio Flows

### v1

@@ 4.2 Sharpening
$\bm{J}_{sharp,i}$ $=-\rho_{i}\Gamma\left[\epsilon\nabla V_{i}-\sum_{i\neq j}V_{i}V_{j}\hat{\bm{n}}_{ij}\right],$ (68)
@@
where $\Gamma$ is the velocity scale parameter, $\epsilon$ is the thickness parameter, $\hat{\bm{n}}_{ij}$ is the pairwise normal, and $V_{ij}$ is the pairwise volume fraction. The second term of Eq. (68) sharpens the interface when two materials are present. Pairs of materials are considered here, so that $\sum_{i=1}^{N_{s}}\bm{J}_{sharp,i}=0$ when the densities of all materials are equal. This prevents diffusion of bulk density from the presence of material interfaces in like materials. The first term of Eq. (68) is an additional diffusion term to prevent over-sharpening. We set $\epsilon=O(\Delta)$, so that the interfaces are not sharpened smaller than the grid scale. $\Gamma$ is a parameter that sets the timescale at which the sharpening term acts. It is assumed that the advection timescale is dominant, so $\Gamma=O(\max(|\bm{u}|))$ to maintain sharp interfaces throughout the simulation. Unless stated otherwise, when sharpening is active, it is assumed that $\epsilon=\Delta$ and $\Gamma=\max(|\bm{u}|)$.
@@ 5.1 Stability Test
First, we repeat the three material stability test introduced in Section 3.1 with the presented method with and without sharpening. Figure 6 shows the stability results for the tradition method and the presented method with and without sharpening. For density ratios of $R>1000$, stability is seen in the presented method with fewer grid points in the interface, demonstrating the stabilizing effects of the formulation. For the largest density ratio cases, such as $R=10^{5}$ and $R=10^{6}$, the sharpening term has a stabilizing effect on the solution relative to the method without sharpening. The pressure error for all cases, with the presented method with and without sharpening is $O(10^{-7}-10^{-12})$ which is 5 to 10 orders of magnitude lower than the traditional method. This stark difference is because the presented scheme maintains equilibrium, while the traditional method does not, as discussed in Sec. 4.1.
@@ 5.2 Shock-Bubble Interaction
Figures 7 show the results at different snapshots in time for both the traditional method and the presented method. Using the presented method, sharp interfaces of uniform thickness are seen throughout the simulation. As the jet reaches the far side of the bubble, a thin filament forms. For the traditional method, this filament persists throughout the simulation as a region of low volume fraction helium. For the presented method, as the filament gets thin, it breaks into small helium bubbles, which later rejoin the larger bubbles on the top and bottom. The sharpening term causes these bubbles to form, because it sharpens the interfaces in mixed regions. The sharpening term sharpens to a thickness based on the $\epsilon$ parameter from Eq. (68) which is a function of the grid size. Hence, the size of these bubbles is based on the numerical grid size, not any physical properties. Experimentally, it was observed in this simulation and others that the sharpening term creates bubbles of radius of approximately $5\Delta$. Hence, the size of the bubbles can be reduced with mesh refinement. For this test case, the density ratio is only $7.25$, so the improved stability from the presented method is not necessary. Additionally, for the case of helium and air, the gases diffuse into one another, so a sharp interface may not be the best way to model these materials. This test case serves to show the effect of the sharp interface on a well-studied test case.
@@
In order to highlight the benefits of the presented method for large density ratio flows, we repeated the shock-bubble interaction case, but decreased the density of helium by $5$ orders of magnitude, so the material properties are $\rho_{1}=1.0$, $\gamma_{1}=1.4$, and $M_{1}=1.0$ and $\rho_{2}=1.38\times 10^{-6}$, $\gamma_{2}=1.67$, and $M_{2}=1.38\times 10^{-6}$. For this case, the density ratio is about $725000$. The artificial diffusivity coefficients are $C_{\mu}=10^{-3}$, $C_{\beta}=0.7$, $C_{D,Y}=C_{D,V}=10^{-2}$, and $\epsilon=1.2\Delta$. As expected from the stability test, the traditional method was unstable for this large of a density ratio while the presented method is able to run stably. Increasing the diffusivity coefficients and decreasing the CFL number were unable to stabilize the traditional method. Figure 8 shows the results for the presented method with and without sharpening at various times. With such a low density in the bubble, the shock quickly passes through it and jetting causes the bubble to break into two stretched halves. Once again, when the sharpening term is used, small bubbles form as the jet is breaking through the far side of the bubble. While this case does not have any physical relevance, it shows that the presented method is able to remain stable for large density ratios with and without sharpening in cases with shocks.
Figure 8: Comparisons between the presented method without sharpening (top) and with sharpening enabled (bottom) for a shock in air causing the breakup of a helium bubble with density $\rho=0.148\times 10^{-5}$. This case is unstable for the traditional LAD method.
@@ 5.3 Rayleigh-Taylor Instability
In order to demonstrate the high-density ratio benefits on a more physically relevant problem, we simulated a high-density ratio Rayleigh-Taylor instability. A $2D$ domain of size $[-64,64]\times[-64,64]\,\text{cm}$ with periodic boundaries in $x$ and wall boundaries in $y$ is discretized with $N=1000$ grid points in each direction. A gravitational acceleration of $-980\,\text{cm}/\text{s}^{2}$ is applied in the $y$ direction. A heavy fluid is initialized above a light fluid with a randomly perturbed interface centered around $y=0\text{cm}$ defined as
@@ 5.4 Droplets in Shear Flow
The simulations were performed on unform grids in both directions with $N=125$, $250$, $500$, and $1000$ grid points in each direction. The initial droplets are defined with profile:
@@ 5.5 Experimental Comparison: Nitrogen Bubble Collapse in Water
[…]
@@
where $z_{0}=0.541$cm and $s(z)$ is the shock waveform. The maximum is taken in the definition of the pressure to avoid negative pressures, which lead to instabilities in mixed points. The shock is defined as planar and its exponential decay is approximated by a Friedlander waveform [54]:
@@
The artificial diffusivity coefficients used are $C_{\mu}=10^{-2}$, $C_{\beta}=1$, $C_{\kappa}=5.0\times 10^{-2}$, $C_{D,Y}=C_{D,V}=1.2$, $\epsilon=1.5\Delta$, and $\Gamma=0.5\max(|\bm{u}|)$. The simulation was run to a final time of $t=1.2\times 10^{-5}\,\text{s}$. The mesh is a uniform grid of $1920\times 3840$ points, which corresponds to $300$ grid points per diameter.
@@
Figure 13: Comparisons between numerical Schlieren computational images (top) and X-ray phase contrast experimental images (bottom) [51] throughout the bubble collapse. Results are plotted at times $t=-0.06,0.14,0.34,0.54,0.74,0.94\,\mu\text{s}$. $t$ is defined such that $t=0\,\mu\text{s}$ is when the shock impacts the bubble. The scale bar in the bottom left corner indicates $100\,\mu\text{m}$.

### v2

@@ 4.2 Sharpening
$\bm{J}^{*}_{sharp,i}$ $=-\rho_{i}\Gamma\left[\epsilon\nabla V_{i}-\sum_{i\neq j}V_{i}V_{j}\hat{\bm{n}}_{ij}\right],$ (68)
@@
where $\Gamma$ is the velocity scale parameter, $\epsilon$ is the thickness parameter, $\hat{\bm{n}}_{ij}$ is the pairwise normal, and $V_{ij}$ is the pairwise volume fraction. The second term of Eq. (68) sharpens the interface when two materials are present. Pairs of materials are considered here, so that $\sum_{i=1}^{N_{s}}\bm{J}^{*}_{sharp,i}=0$ when the densities of all materials are equal. This prevents diffusion of bulk density from the presence of material interfaces in like materials. The first term of Eq. (68) is an additional diffusion term to prevent over-sharpening. We set $\epsilon=O(\Delta)$, so that the interfaces are not sharpened smaller than the grid scale. $\Gamma$ is a parameter that sets the timescale at which the sharpening term acts. It is assumed that the advection timescale is dominant, so $\Gamma=O(\max(|\bm{u}|))$ to maintain sharp interfaces throughout the simulation. Unless stated otherwise, when sharpening is active, it is assumed that $\epsilon=\Delta$ and $\Gamma=\max(|\bm{u}|)$.
@@
4.3 Avoiding Unphysical Bulk Density Diffusion
In the presented method, artificial bulk density diffusion is a mechanism by which interfaces are regularized. The diffusion term and the sharpening term are designed to ensure they do not cause bulk density change at interfaces between materials with equal densities. This can be shown as follows:
Assume the domain contains $N_{s}$ materials of densities $\rho_{i}=\rho_{0}$. Summing Eq. (1) across all materials gives:
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\bm{u})$ $=-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\bm{J}^{*}_{i}\right)$ (72)
$\frac{\partial\rho}{\partial t}+\rho_{0}\nabla\cdot(\bm{u})$ $=-\nabla\cdot\left(\sum_{i=1}^{N_{s}}\left(-D_{\rho}^{*}\nabla(\rho Y_{i})-\rho_{i}\Gamma\left[\epsilon\nabla V_{i}-\sum_{i\neq j}V_{i}V_{j}\hat{\bm{n}}_{ij}\right]\right)\right)$ (73)
$=\nabla\cdot(D_{\rho}^{*}\nabla\rho_{0})+\nabla\cdot\left(\rho_{0}\Gamma\left[\nabla\left(\sum_{i=1}^{N_{s}}V_{i}\right)-\sum_{i=1}^{N_{s}}\sum_{i\neq j}V_{i}V_{j}\hat{\bm{n}}_{ij}\right]\right)$ (74)
$=\nabla\cdot(D_{\rho}^{*}\nabla\rho_{0})$ (75)
$=0.$ (76)
Hence, the artificial diffusion terms do not cause a change in bulk density if all of the materials have the same density. This is the result of two main choices in the design of the method. First, the artificial species diffusivity for all materials are set to the same value, $D_{\rho}^{*}$. This choice allows the sum over the materials to enter into the gradient of the first term on the right-hand side of Eq. (73) and ultimately go to zero. In Eq. (24), the traditional method computes a separate material diffusivities based on ringing and overshoots of each material’s mass fraction. For the case of two materials, $\overline{F(Y_{l})}=\overline{F(Y_{h})}$ and $\overline{|Y_{l}|-1+|1-Y_{l}|}=\overline{|Y_{h}|-1+|1-Y_{h}|}$, so with the traditional method, $D^{*}_{l}=D^{*}_{h}$, which satisfied the desired condition. However, for more than two materials, there are no such guarantees for the traditional diffusivity, so it could cause diffusion of density in a constant density interface. Hence, the choice of a common diffusivity across all materials is crucial to the formulation for more than two materials.
The second design choice is the use of the N-phase CDI formulation from Mirjalili [35] for the sharpening. Considering the materials pairwise while sharpening ensures that the sharpening terms sum to zero when summed across all materials. Without such a formulation, the density of constant density flows could diffuse from the sharpening term. Section 5.4 shows a three material, constant density case that demonstrates this strength of the presented method.
4.4 Comparison with Artificial Mass Diffusion Methods
In this section, we discuss the differences between the presented method and other LAD methods using artificial mass diffusion in the literature. The work of Terashima [54] uses mass diffusion for stabilization of multi-material interfaces without sharpening. To enforce interface equilibrium, they restrict their materials to calorically perfect gases and introduce a PDE to solve for the mixture ratio of specific heats. Aslani [19] introduced a five-equation model for stiffened gases with artificial mass diffusion with no sharpening. They achieve interface equilibrium through a similar formulation of consistent fluxes as presented, but formulate the energy equation for the stiffened gas equation of state. The work of Jain [29] presents a five-equation model with a similar diffusion and sharpening formulation for two material flows. Instead of formulating the artificial diffusion fluxes in a consistent way, they use carefully constructed numerical fluxes to ensure interface equilibrium is preserved. The work of Collis [47] presents a four-equation model for multi-material and multi-phase flows using a CDI formulation for sharpening. This formulation uses similar consistent flux terms, but enthalpy is used in the energy consistency term, as opposed to internal energy in the presented formulation. Additionally, it is restricted to stiffened-gas equations of state and only tests two-material cases. The presented method expands upon the existing literature by ensuring interface equilibrium is maintained in a four-equation model for any number of materials without any assumptions on the equations of state. This is primarily achieved through the choices of a single artificial bulk density diffusivity, $D_{\rho}^{*}$, across all materials, careful derivation of the consistency requirements in Eqs. (38)-(40), and the formulation of the sharpening term considering pairs of materials.
@@ 5.1 Stability Test
First, we repeat the three material stability test introduced in Section 3.1 with the presented method with and without sharpening. Figure 6 shows the stability results for the tradition method and the presented method with and without sharpening. For density ratios of $R\geq 100$, stability is seen in the presented method with fewer grid points in the interface, demonstrating the stabilizing effects of the formulation. Figure 6 also shows that the sharpening term does not affect the stability regimes of the presented LAD method. The pressure error for all cases, with the presented method with and without sharpening is $O(10^{-6}-10^{-12})$ which is 5 to 10 orders of magnitude lower than the traditional method. This stark difference is because the presented scheme maintains equilibrium, while the traditional method does not, as discussed in Sec. 4.1.
@@ 5.2 Shock-Bubble Interaction
For this test case, the density ratio is only $7.25$, so the improved stability from the presented method is not necessary. Additionally, for the case of helium and air, the gases diffuse into one another, so a sharp interface may not be the best way to model these materials. Instead, this test case serves to demonstrate the effect of the interface sharpening on a well-studied test case. Figure 7 shows the results at different snapshots in time for both the traditional method and the presented method. Using the presented method, sharp interfaces of uniform thickness are seen throughout the simulation. As the jet reaches the far side of the bubble, a thin filament forms. For the traditional method, this filament persists throughout the simulation as a region of low volume fraction helium. For the presented method, as the filament gets thin, it breaks into small helium bubbles, which later rejoin the larger bubbles on the top and bottom. The sharpening term causes these bubbles to form, because it sharpens the interfaces in mixed regions. The sharpening term sharpens to a thickness based on the $\epsilon$ parameter from Eq. (68) which is a function of the grid size. Hence, the size of these bubbles is based on the numerical grid size, not any physical properties. Experimentally, it was observed in this simulation and others that the sharpening term creates bubbles of radius of approximately $5\Delta$. Hence, the size of the bubbles can be reduced with mesh refinement. In this case, the sharp interface of the presented method is not expected to be more representative of the true physics, but this case serves as a good demonstration of the properties that the method displays.
@@
In order to highlight the benefits of the presented method for large density ratio flows, we repeated the shock-bubble interaction case, but decreased the density of helium by $5$ orders of magnitude, so the material properties are $\rho_{1}=1.0$, $\gamma_{1}=1.4$, and $M_{1}=1.0$ and $\rho_{2}=1.38\times 10^{-6}$, $\gamma_{2}=1.67$, and $M_{2}=1.38\times 10^{-6}$. For this case, the density ratio is about $725000$. The artificial diffusivity coefficients are $C_{\mu}=10^{-3}$, $C_{\beta}=0.7$, $C_{D,Y}=C_{D,V}=10^{-2}$, and $\epsilon=1.2\Delta$. This larger value of $\epsilon$ was observed to better stabilized this large density ratio in the prescence of the shock and had minimal effect on the thickness of the interface. As expected from the stability test, the traditional method was unstable for this large of a density ratio while the presented method is able to run stably. Increasing the diffusivity coefficients and decreasing the CFL number were unable to stabilize the traditional method. Figure 8 shows the results for the presented method with and without sharpening at various times. With such a low density in the bubble, the shock quickly passes through it and jetting causes the bubble to break into two stretched halves. Once again, when the sharpening term is used, small bubbles form as the jet is breaking through the far side of the bubble. While this case does not have any physical relevance, it shows that the presented method is able to remain stable for large density ratios with and without sharpening in cases with shocks.
Figure 8: Comparisons between the presented method without sharpening (top) and with sharpening enabled (bottom) for a shock in air causing the breakup of a helium bubble with density $\rho=0.138\times 10^{-5}$. This case is unstable for the traditional LAD method.
@@ 5.3 Rayleigh-Taylor Instability
In order to demonstrate the high-density ratio benefits on a more physically relevant problem, we simulated a high-density ratio Rayleigh-Taylor instability. A $2D$ domain of size $64,64]\times[-64,64]\,\text{cm}$ with periodic boundaries in $x$ and wall boundaries in $y$ is discretized with $N=1000$ grid points in each direction. A gravitational acceleration of $980\,\text{cm}/\text{s}^{2}$ is applied in the $y$ direction. A heavy fluid is initialized above a light fluid with a randomly perturbed interface centered around $y=0\,\text{cm}$ defined as
@@ 5.4 Droplets in Shear Flow
The simulations were performed on uniform grids in both directions with $N=125$, $250$, $500$, and $1000$ grid points in each direction. The initial droplets are defined with profile:
@@ 5.5 Experimental Comparison: Nitrogen Bubble Collapse in Water
Finally, we compare results obtained with the presented model to experimental recordings of the shock-induced dynamics of a spherical nitrogen bubble in water [57]. In brief, a microfluidic device generates spherical nitrogen bubbles in a water-filled test chamber. A flat copper foil is driven to electrical explosion $6\,\text{mm}$ away from the bubble with a high energy pulsed power driver, which induces a planar shock wave [58]. The shock wave propagation and bubble dynamics resulting from their interaction are recorded through in-situ ultra-high-speed synchrotron-based X-ray phase contrast imaging at $5\,\text{Mfps}$. The peak pressure of the shock can be deduced from its observable speed and the Rankine-Hugoniot jump relation and its temporal decay is estimated through hydrodynamic numerical simulations [59]. More information on these experimental recordings are available at [57].
@@
where $z_{0}=0.541\,\text{cm}$ and $s(z)$ is the shock waveform. The maximum is taken in the definition of the pressure to avoid negative pressures, which lead to instabilities in mixed points. The shock is defined as planar and its exponential decay is approximated by a Friedlander waveform [60]:
@@
The artificial diffusivity coefficients used are $C_{\mu}=10^{-2}$, $C_{\beta}=1$, $C_{\kappa}=5.0\times 10^{-2}$, $C_{D,Y}=C_{D,V}=1.2$, $\epsilon=1.5\Delta$, and $\Gamma=0.5\max(|\bm{u}|)$. These coefficients were found to be the minimum robust values to stabilize the pressure pulse’s impact with the bubble. The simulation was run to a final time of $t=1.2\times 10^{-5}\,\text{s}$. The mesh is a uniform grid of $1920\times 3840$ points, which corresponds to $300$ grid points per diameter.
@@
Figure 13: Comparisons between numerical Schlieren computational images (top) and X-ray phase contrast experimental images (bottom) [57] throughout the bubble collapse. Results are plotted at times $0.06,0.14,0.34,0.54,0.74,0.94\,\mu\text{s}. tt$ is defined such that $t=0\,\mu\text{s}$ is when the shock impacts the bubble. The scale bar in the bottom left corner indicates $100\,\mu\text{m}$.